# TLControl: Trajectory and Language Control for Human Motion Synthesis

### v3

@@ 4 Experiment
Datasets
Method Control Joint FID $\downarrow$ R-precision $\uparrow$ (Top-3) Diversity $\rightarrow$ Traj. Err. $\downarrow$ (50 cm, %) Loc. Err. $\downarrow$ (50 cm, %) Avg. Err. (cm) $\downarrow$
@@
Metrics
Following the experiments setting in [54], we evaluate the fidelity of our result using Frechet Inception Distance (FID), R-Precision and Diversity. FID measures the quality of motion created by generative techniques. It calculates the disparity between actual and produced distributions within the feature space of a pre-trained model as being used in [11]. R-Precision measures a motion sequence against 32 text narratives, one of which is accurate and the remaining 31 are random descriptions. It involves computing the Euclidean distances between the embeddings of the motion and the textual narratives. The accuracy of retrieving the correct text from the motion is then reported based on the top-3 matches. Diversity is evaluated by randomly pairing all generated sequences from the test texts. The average cumulative difference within each pair is then calculated to determine diversity.
Implementation Details
In the part-based VQ-VAE, the settings of the encoder and the decoder module are based on T2M-GPT [58] with the downsampling $s$ factor set to 4. For the masked trajectory transformer, we use a frozen CLIP-ViT-B/32 [37] model for pre-processing the text prompt $\bm{L}$, and then a 4-layer transformer processes the text information and control trajectories $\bm{R^{\prime}}$ into the latent features, followed by a 3-layer transformer decodes the latent features into the codes indices. The embedding dimension of these two stage transformers are 512 and 256 respectively. The networks are trained using AdamW optimizer with learning rate decayes from 1e-4 to 1e-7 with a batch size of 64. For the run-time optimization step, we use Limited-Memory BFGS [26] as the optimization algorithm with learning rate set to 0.1. The accuracy criteria is set to 1E-6 and we provide an ablation study of the criteria setting in  4.3.4.
In this experiments section, we use a machine with RTX4090 GPU for running the tests except in  4.2 for a fair comparison. The testing batch size is 32 following previous works using HumanML3D [11] and KIT-ML [35] dataset.
Figure 3: Qualitative results of our method. Figure 3a and Figure 3d demonstrate that our method enables separate controls using language and joint-level trajectories. Figure 3b and Figure 3c showcase the capability of our method to manage multi-joint control simultaneously. Please refer to our supplementary for more qualitative results
@@ 4.1 Controllable Human Motion Generation
We compare our method in controllable human motion generation with the current SOTA methods. When comparing with MDM [48], PriorMDM [39], and GMD [22], we focus on controlling the pelvis only for fair comparisons. We further compare with Omnicontrol [54] with more settings of different combinations of control joints. The results are shown in 1 and 2. For the single joint controlling task, our method achieves better scores in R-precision while having comparable results in other fidelity metrics. However, our approach consistently surpasses existing methodologies in the metrics measuring control accuracy. On the Humanml3D test set, the results generated by our approach steadily remained within a 50 cm range of the intended control signals, showcasing remarkable accuracy and stability. We also demonstrate enhanced control accuracy on KIT dataset for using the pelvis only and the average of controlling the joints individually, where we continues to outperform other methods in accuracy.
In the task that utilize all six joints for controlling, Omnicontrol struggles with multi-joint controls. In contrast, our method can handle multi-joint control effectively, producing highly realistic outcomes with such abundant control signals. Visual comparison is shown in Fig  4.
Methods Ours OmniControl MDM GMD
Time (s/frame) normal-$\downarrow$ 0.016 $0.619$ $0.200$ $0.561$
FPS normal-$\uparrow$ 61.7 $1.6$ $5$ $1.8$
Table 3: Runtime Efficiency of different methods.
@@ 4.2 Run-time Performance
We further compared our method with the SOTA methods in terms of run-time performance. We run our method on a machine with an RTX2080Ti GPU and record the average process time for a sample using one control joint. As shown in Tab. 3, due to our structured and compact latent space and efficient optimization framework, our method consistently outperforms all existing methods with a significant margin.
Figure 4: Qualitative comparison results. We compare our results with Omnicontrol[54]. Our results are shown in Yellow, while those of Omnicontrol are depicted in Green. Please refer to our supplementary for more comparison results
@@ 4.3.1 Motion Latent Space
We investigate the learned part-based VQ-VAE for motion embedding. In Fig. 5, we randomly choose 10 test cases from Humanml3D test set, and process them with MTT for achieving the corresponding features. We present the result of using t-SNE [50] for visualizing these features, offering insights into the structure of the acquired motion latent space. Our part-based VQ-VAEcontributes a compact motion latent space while preserving the semantics described by language inputs, as evidenced in Fig. 5.
Figure 5: t-SNE visualization of learned motion latent space. Each color stands for a textural description of the motion. We randomly select ten sentences for visualization.
Figure 6: Motion variations controlled by language (text in blue) and randomly sampled body part level latent codes. Three sample motions in red, blue, and green are shown.
Next, we visualize the learned body part-level motion prior. As shown in Fig. 6, our framework for motion embedding learns a well-structured latent space. This results in generating motions that conform to the language description while capturing a diverse distribution of motions under each text condition.
4.3.2 Trajectory Incompleteness
We investigate the performance of our method under different levels of trajectory incompleteness, for which we simulate various incompleteness scenarios by adjusting the ratio of random masking. As revealed in Fig. 7, our model’s performance degrades gracefully even as the control trajectories become more incomplete. As a reference, note that even with full joint trajectories, the performance of previous methods [54] in terms of FID and R-precision are $2.614$ and $0.606$.
@@ 4.3.2 Trajectory Incompleteness
Table 4: Average Error for Different Joints v.s. Accuracy criteria during optimization when controlling all the joints.
4.3.3 Runtime of different joint control strategy
We present the runtime associated with different joint control strategies. In Figure  8, we record the average runtime for processing a batch under the condition of using different combinations of joint trajectories. Due to the complexity of hand motions, optimization that exclusively using both hands trajectories is the most time-consuming.
@@ 5 Conclusion
We introduce TLControl, a method for controllable human motion generation using a combination of joint trajectory and language inputs. TLControl leverages a disentangled latent space for diverse human motion, allowing for effective optimization and producing high-fidelity motion aligned with both language descriptions and specified trajectories. The proposed Masked Trajectories Transformer captures the correspondence between trajectories and textual information, leading to a well-structured compact latent space. We further introduce an efficient optimization framework for joint-level motion editing while maintaining motion semantics encoded by the language, thus enabling interactive generation and modification of high-quality motion within a short run-time. Extensive experiments validate the effectiveness of our framework, highlighting its capability to enable users to quickly generate and modify high-fidelity human motion interactively.
6 Acknowledgements
The authors would like to thank Yiming Xie for preparing sample cases in the comparisons. Dinesh Jayaraman was supported by NSF CAREER Award 2239301.
@@ 6 Acknowledgements
This supplementary material covers: the abalation study of our part-based VQ-VQE embedding (Sec. 1); an experiments of our multi-joint control results (Sec. 2); more qualitative comparisons with the SOTA methods (Sec. 3); more implementation details (Sec. 4); and a discussion of limitations and future work (Sec. 5).
@@
1 Ablation Study on Part-based VQ-VAE
@@
To validate the effectiveness of the proposed part-based VQ-VAE, we compared our network structure with a counterpart VQ-VAE without body-part level disentangling, which we referred to as Unsplit VQ-VAE. To guarantee a fair comparison, both models were configured with an identical feature dimension that represents the full-body motion and is utilized for the runtime optimization. As summarized in Tab. 1, our proposed part-based VQ-VAE demonstrated superior performance in generating controllable motion, showing its efficacy in preserving language semantics alongside facilitating flexible trajectory control. This is attributed to our disentanglement design of our VQ-VAE that allows an efficient representation of motion dynamics at the body-part level. Moreover, our part-based VQ-VAE exhibits the capability to encode motions into a more compact latent space. This compactness not only enhances the model’s accuracy in handling complex motion but also significantly reduces processing time during the runtime optimization, making it more suitable for real-time applications where rapid motion generation and adaptation are crucial.
@@
Figure 2: Qualitative comparison results with GMD [22] in the task of controlling root path. Our results are shown in Yellow, while the results of GMD are depicted in Green. Please refer to our supplementary video for more details of the comparison.
2 Multi-Joint Control Results
@@
In this section, we present the performance metrics of our method when applied to various groups of joints. This test is performed in the test set of HumanML3D dataset [11], and the results are shown in Tab. 2. Our approach demonstrates notable efficiency and accuracy in managing multiple joints simultaneously. As we integrate additional control trajectories, the motion representation becomes increasingly detailed. This enhancement allows our method to generate more realistic results, showcasing its adaptability and precision in complex joint operations.
3 More Qualitative Comparisons on Controllable Motion Synthesis
We present more detailed qualitative results in this section, comparing our method with the SOTA methods in terms of controllable human motion synthesis: GMD [22] and Omnicontrol [54].
Given that GMD only concentrates on the trajectory control of the root joint, we demonstrate cases where only the root path is assigned, as depicted in Figure 2. Omnicontrol allows control over multiple joints simultaneously. To illustrate this, we compare tasks that involve controlling different joints concurrently with Omnicontrol, with the results displayed in Figure 3. Our method demonstrates superior accuracy in adhering to the control trajectories and achieves this with a notably reduced runtime. Our approach exhibits enhanced adaptability and precision in complex motion patterns, outperforming the compared methods in dynamic motion scenarios where precise joint control is critical.
4 Implementation Details
@@
In implementing the masked trajectory transformer, we employ a frozen CLIP-ViT-B/32 [37] model for the initial processing of the text prompt, which translates it into a 512-dimensional vector representative of language features. The joint trajectories are grouped with every 4 closest waypoints being bundled into a single temporal token. This bundling is reflective of our choice to use a downsampling rate of 4 within the part-based VQ-VAE framework. Each waypoint group is then mapped onto a 512-dimensional feature space. A standard 4-layer transformer encoder is used to integrate the language and trajectory features into a cohesive latent space, where tokens comprise both the language feature vector and the feature representations of the trajectory at each temporal token. Following this, a 3-layer transformer encoder is designed to transform the latent space into a sequence of logits for determining the code indices. To transform logits into selected code indices, we follow the common idea of applying softmax and then argmax, but in a differentiable manner: we use the Gumbel-Softmax [^1] Gumbel-Softmax:  https://pytorch.org/docs/stable/generated/torch.nn.functional.gumbel_softmax.html for producing a one-hot decision vector $\mathbf{v}$, followed by selecting the most probable code index using $\mathbf{C}_{k}\cdot\mathbf{v}$ where $\mathbf{C}_{k}\in\mathbb{R}^{d\times|\mathbf{C}|}$ is the codebook for a joint group containing $|\mathbf{C}|$ distinct codes, each of dimension $d$. Then the selected codes are concatenated to serve as the initial full-body state for the VQ-decoder.
@@
5 Limitations and Future Work
Although TLControl achieves controllable motion synthesis that is faithfully aligned with language semantics and control trajectories with high efficiency, there are still some limitations. In particular, the absence of a physics-based simulation in the current version of TLControl may produce human motions that deviate from established physics rules. Thus, one promising direction is to incorporate physics-based simulation [29, 8, 30, 27] into the generation process, as evidenced by [56, 20].
Furthermore, our approach facilitates detailed motion synthesis, enabling its application in the generation of motions responsive to scenes or objects. This can be achieved by defining control trajectories that align with the movements of specific objects or scenes. However, more effort can be made to realize scene-aware motion synthesis that accurately responds to dynamic changes, e.g., simulating walking on terrains with diverse elevations [16, 52, 53, 43, 57], which can be an intriguing avenue for future research.

### v4

@@ 4.0.2 Metrics
4.0.3 Implementation Details
In the part-based VQ-VAE, the settings of the encoder and the decoder module are based on T2M-GPT [67] with the downsampling $s$ factor set to 4. For the masked trajectory transformer, we use a frozen CLIP-ViT-B/32 [43] model for pre-processing the text prompt $\bm{L}$, and then a 4-layer transformer processes the text information and control trajectories $\bm{R^{\prime}}$ into the latent features, followed by a 3-layer transformer decodes the latent features into the codes indices. The embedding dimension of these two stage transformers are 512 and 256 respectively. The networks are trained using AdamW optimizer with learning rate decays from 1e-4 to 1e-7 with a batch size of 64. For the run-time optimization step, we use Limited-Memory BFGS [32] as the optimization algorithm with learning rate set to 0.1. The accuracy criteria is set to 1E-6 and we provide an ablation study of the criteria setting in Sec. 4.3.2.
In this experiments section, we use a machine with RTX4090 GPU for running the tests except in Sec. 4.2 for a fair comparison. The testing batch size is 32 following previous works using HumanML3D [15] and KIT-ML [41] dataset.
@@ 4.1 Controllable Human Motion Generation
4.1.1 Qualitative Results
In Fig. 3 and Fig. 4, we showcase some qualitative results of our method using various control trajectories from: 1) Specifying parameters of analytical shapes; 2) Hand drawings in 3D space; and 3) Environment settings. This versatility underlines our method’s practical applicability across a wide range of scenarios, demonstrating its significant utility in diverse applications.
Root Traj Masked Rate 0% 25% 50% 75%
MModality $\uparrow$ 0.994 1.379 1.790 1.982
Table 3: The MModality of the generated motions under different trajectory masking rates.
4.1.2 Diverse Output Sampling
Our approach enables the sampling of multiple outputs for the same trajectory and language input, as illustrated in Figure 6. We conduct an experiment on the test set of HumanML3D, showing the sampling capability of our method by using text and root trajectory in Tab. 3.
Figure 6: Qualitative results under the same control trajectories and text prompts.
Figure 7: Qualitative comparison with Omnicontrol[62] using the samples from the test set of HumanML3D [15]. Our results are in Yellow and Omnicontrol’s in Green.
Figure 8: Qualitative comparison results with GMD [26] in the task of controlling root path. Our results are in Yellow and GMD’s in Green.
4.1.3 Comparison with SOTAs
We compare TLControl in controllable motion generation with the SOTA methods. When comparing with MDM [56], PriorMDM [45], and GMD [26], we focus on controlling the pelvis only for fair comparisons. We further compare with Omnicontrol [62] with more settings of using different end-effectors and the combination of joints. The results are shown in Tab. 1 and Tab. 2. For the single joint controlling task, our method achieves better scores in R-precision while having comparable results in other fidelity metrics. Meanwhile, our method consistently surpasses other models in the metrics measuring control accuracy. On the HumanML3D, the results generated by our method steadily remained within a 50 cm range of the intended control signals, showcasing remarkable accuracy and stability. We also demonstrate enhanced control accuracy on KIT dataset for using the pelvis only and the average of controlling the joints individually, where TLControl consistently outperforms other methods in accuracy. When using all six joints for controlling, Omnicontrol struggles with multi-joint controls. In contrast, our method can handle multi-joint control effectively, producing highly realistic outcomes with such abundant control signals. Visual comparison is shown in Fig. 5, Fig. 8 and Fig. 8.
Methods Time (s/frame) $\downarrow$
OmniControl 0.606
MDM 0.192
GMD 0.567
Ours 0.015
Table 4: Runtime of different methods.
@@ 4.2 Run-time Performance
We further compared our method with the SOTA methods in terms of run-time performance. We run the methods on a machine with an RTX3080Ti GPU and record the average process time for a test sample from HumanML3D [15]. As shown in Tab. 4, due to our structured and compact latent space and efficient optimization framework, our method consistently outperforms all existing methods by a significant margin.
Figure 9: t-SNE visualization of learned latent space. Each color stands for a textural description of the motion. We randomly select ten sentences for visualization.
Figure 10: Motion diversity controlled by language (blue text) and randomly sampled body part latent codes. Three sample motions are in red, blue, and green.
@@ 4.3.1 Motion Latent Space
We investigate the learned part-based VQ-VAE for motion embedding. In Fig. 10, we randomly choose 10 test cases from HumanML3D test set, and process them with MTT for achieving the corresponding features. We present the result of using t-SNE [34] for visualizing these features, offering insights into the structure of the acquired motion latent space. Our part-based VQ-VAE contributes a compact motion latent space while preserving the semantics described by language inputs, as evidenced in Fig. 10. Next, we visualize the learned body part-level motion prior. As shown in Fig. 10, our framework for motion embedding learns a well-structured latent space. This results in generating motions that conform to the language description while capturing a diverse distribution of motions under each text condition.
@@
Table 5: Average Joint Error v.s. Accuracy criteria during optimization when controlling all the joints.
@@ 5 Conclusion
We introduce TLControl, a method for controllable human motion generation using a combination of joint trajectory and language inputs. TLControl leverages a human body morphology-aware structured latent space using a VQ-VAE, and a motion distribution conditioned on language and trajectory using a Masked Trajectory Transformer. We further introduce an efficient optimization framework offers flexible controls by allowing users to specify various optimization goals while ensuring high runtime efficiency. Extensive experiments validate the effectiveness of our framework, highlighting its capability to enable users to quickly generate high-fidelity human motion interactively.
For the cases where there exist conflicts between input trajectories and languages, generating motions under these conflicting inputs is an open question. Handling such cases is beyond the scope of this paper. In this paper, we assume that the language and trajectory controls do not conflict. Nevertheless, we acknowledge that this issue presents a compelling avenue for future research. Future enhancements could focus on conflict detection and intuitive user guidance for adjustments for more accurately aligned animations with user expectations.
Acknowledgment
This work is partly supported by the Innovation and Technology Commission of the HKSAR Government under the ITSP-Platform grant (Ref: ITS/319/21FP) and the InnoHK initiative (TransGP project). The authors would like to thank Yiming Xie for preparing sample cases in the comparisons. Dinesh Jayaraman was supported by NSF CAREER Award 2239301.
@@ Acknowledgment
This supplementary material covers: the ablation study of our part-based VQ-VQE embedding (Sec. 0.A); an experiments of our multi-joint control results (Sec. 0.B); the performance of our method using incompleted trajecotries (Sec. 0.C); the runtime of different joint control strategy (Sec. 0.D); an ablation study with IK based solution (Sec. 0.E); results of using only one modality input (Sec. 0.F); more implementation details (Sec. 0.G); and more qualitative results (Sec. 0.H);.
@@
Appendix 0.A Ablation Study on Part-based VQ-VAE
@@
To validate the effectiveness of the proposed part-based VQ-VAE, we compared our network structure with a counterpart VQ-VAE without body-part level disentangling, which we referred to as Unsplit VQ-VAE. To guarantee a fair comparison, both models were configured with an identical feature dimension that represents the full-body motion and is utilized for the runtime optimization. As summarized in Table 1, our proposed part-based VQ-VAE demonstrated superior performance in generating controllable motion, showing its efficacy in preserving language semantics alongside facilitating flexible trajectory control. This is attributed to our disentanglement design of our VQ-VAE that allows an efficient representation of motion dynamics at the body-part level. Moreover, our part-based VQ-VAE exhibits the capability to encode motions into a more compact latent space. This compactness not only enhances the model’s accuracy in handling complex motion but also significantly reduces processing time during the runtime optimization, making it more suitable for real-time applications where rapid motion generation and adaptation are crucial.
@@
Appendix 0.B Multi-Joint Control Results
@@
In this section, we present the performance metrics of our method when applied to various groups of joints. This test is performed in the test set of HumanML3D dataset [15], and the results are shown in Table  2. Our approach demonstrates notable efficiency and accuracy in managing multiple joints simultaneously. As we integrate additional control trajectories, the motion representation becomes increasingly detailed. This enhancement allows our method to generate more realistic results, showcasing its adaptability and precision in complex joint operations.
Appendix 0.C Trajectory Incompleteness
We investigate the performance of our method under different levels of trajectory incompleteness, for which we simulate various incompleteness scenarios by adjusting the ratio of random masking. As revealed in Figure  2, our model’s performance degrades gracefully even as the control trajectories become more incomplete. As a reference, note that even with full joint trajectories, the performance of previous methods [62] in terms of FID and R-precision are $2.614$ and $0.606$.
Appendix 0.D Runtime of Different Joint Control Strategy
We present the runtime associated with different joint control strategies. In Figure  3, we record the average runtime for processing a batch under the condition of using different combinations of joint trajectories. Due to the complexity of hand motions, optimization that exclusively using both hands trajectories is the most time-consuming.
Appendix 0.E Ablation with IK Based Solution
Control Strategy No-opt Joint-IK Ours (Latent-Opt)
Traj Masked Rate 0% 50% 0% 50% 0% 50%
FID (Root) $\downarrow$ 0.343 0.391 0.336 0.675 0.271 0.311
FID (LHand) $\downarrow$ 0.226 0.297 0.383 0.612 0.135 0.224
Table 3: Quantitative results of comparing IK based solution.
Furthermore, we conduct a comparative analysis between the use of inverse kinematics (IK) and our optimization framework. In this experiment, we control the root and left hand separately using trajectories with different mask rates. We report the FID for motions: 1) decoded from initial state features without optimization; 2) optimized from the decoded motions over joints using IK; 3) optimized from the decoded motions over the latent features using our design. Spatially, IK is for joint level adjustment, and it is limited in maintaining the semantics of the overall body motion. Temporally, in cases of incomplete control trajectories, IK cannot guide the frames without trajectories, leading to discrepancies between the generated motions and the intended control goals.
Appendix 0.F Using Only One Modality Input
Figure 4: Qualitative results of using only one modality.
TLControl is compatible with empty trajectories $R$ or empty language input $L$ for pure language-conditioned or trajectories-conditioned motion generation, although we typically assume that trajectory and language descriptions are both given. Here we give examples of using language only and using trajectories only for motion generation in Figure 4, where our method produces high-quality results.
Appendix 0.G Implementation Details
@@
In implementing the masked trajectory transformer, we employ a frozen CLIP-ViT-B/32 [43] model for the initial processing of the text prompt, which translates it into a 512-dimensional vector representative of language features. The joint trajectories are grouped with every 4 closest waypoints being bundled into a single temporal token. This bundling is reflective of our choice to use a downsampling rate of 4 within the part-based VQ-VAE framework. Each waypoint group is then mapped onto a 512-dimensional feature space. A standard 4-layer transformer encoder is used to integrate the language and trajectory features into a cohesive latent space, where tokens comprise both the language feature vector and the feature representations of the trajectory at each temporal token. Following this, a 3-layer transformer encoder is designed to transform the latent space into a sequence of logits for determining the code indices. To transform logits into selected code indices, we follow the common idea of applying softmax and then argmax, but in a differentiable manner: we use the Gumbel-Softmax [23] for producing a one-hot decision vector $\mathbf{v}$, followed by selecting the most probable code index using $\mathbf{C}_{k}\cdot\mathbf{v}$ where $\mathbf{C}_{k}\in\mathbb{R}^{d\times|\mathbf{C}|}$ is the codebook for a joint group containing $|\mathbf{C}|$ distinct codes, each of dimension $d$. Then the selected codes are concatenated to serve as the initial full-body state for the VQ-decoder.
@@
Appendix 0.H More Qualitative Comparisons on Controllable Motion Synthesis
Figure 5: Qualitative results of our method. Please refer to our video for more results.
In Figure 5, we showcase additional qualitative results. For more demos and comparisons with other methods, please refer to our supplementary video.
Our method not only demonstrates superior accuracy in adhering to control trajectories and achieves this with a notably reduced runtime but also excels in generating motions that accurately reflect the semantics of text prompts. Our approach exhibits enhanced adaptability and precision in complex motion patterns, outperforming compared methods in dynamic motion scenarios where precise joint control is critical and the ability to capture the essence of textual instructions is paramount.